A 10% solution is mixed with 1L of water to get 5L of a 4% solution. how much of the 10% solution was available at the start?
Answer by stanbon(75887)

A 10% solution is mixed with 1L of water to get 5L of a 4% solution. how much of the 10% solution was available at the start?
----
Equation:
water + water = water
0.90x + 1*1 = 0.96(5)
----
90x + 100 = 96*5
90x = 380
x = 3.22 liters (amt. of 10% solution at the start)
===================
Cheers,
Stan H.
====================